I got this as homework, i'm in grade 5, and I just need some help.
A1.Express Each sentence as a mathematical equation.
a)five times a number is equal to fifteen.
b)Twice Mary's age is thirty.
c)A number increased by seven is equal to thirteen.
d)A number decreased by two is equal to eleven.
e)Six less that a number is equal to fourteen.
f)Bill's age two years ago was ten.
g)Nicole's age five years from now will be twenty-one.

A2.What is the coeffecient of each term?
a)5w
b)-3w
c)-7y
d)m
e)-4.3m
f)-6
g)-m
h)3/4m
A3.Solving
a)x^5*x^7=
b)a^9/a^5=
c)y^12

Answer by heidi(17)

a. 5 x n = 15 (mary)
b. 2 x 15 = 30
c.n + 7 = 13 (n- 6)
d.n - 2 = 11 (13)
e.6 - n = 14 (20)
f. n = a + 2
g. n = 21 -5
how could u be doing the rest of this in 5th grade ?